OpenStudy (itsbribro):

Please Help!!!! #1.) What is the point-slope form of the equation of a line that passes through the point (1, 4) and has a slope of -3? A. y - 1 = -3 (x-4) B. y + 4 = 3 (x+1) C. y + 4 = -3 (x+1) D. y - 4 = -3 (x-1)

OpenStudy (anonymous):

Do you need solution or just answer?

OpenStudy (itsbribro):

#2.) A yearbook printer charges based on the number of pages printed. Here is a table that shows the cost of some recent yearbooks: http://static.k12.com/calms_media/media/241500_242000/241736/1/ba4f4a8c374db4825eb2e52a5d38ac0ace85d489/07_08_LQ_4.gif Which equation gives the cost, y, in terms of the number of pages, x? A. y = 0.012x + 0.65 B. y = 83.3x + 0.65 C. y = 0.012x + 1.25 D. y = 0.012x + 49.985

OpenStudy (anonymous):

C for the first question

OpenStudy (anonymous):

A for the second one

OpenStudy (itsbribro):

i think its A

OpenStudy (itsbribro):

lol thanks